OpenStudy (dls):

find the 8th term of sequence 2/11,1/5,2/9

OpenStudy (anonymous):

should be just find the common ratio and use the geometric sequence formula

OpenStudy (dls):

Is it even in GP?

OpenStudy (anonymous):

gp?

OpenStudy (dls):

geometric progression

OpenStudy (anonymous):

divide and see

OpenStudy (anonymous):

but i got no it isnt

OpenStudy (dls):

i now it isnt

OpenStudy (anonymous):

This sequence does not represent any conventional series. Can you revise that question and/or give us any other details about the question. For example, does it have a part (a) ?

OpenStudy (dls):

No this is complete

OpenStudy (anonymous):

i forgot what you write for the answer but i assume it would be "not a sequence" or something like that

hartnn (hartnn):

isn't the series in HP?? 11/2+9/2=10 2/(1/5)=10 so just use formula for n th term of HP

OpenStudy (anonymous):

eigth term = 2/3